: Ch 4 20 of 20 (16 complete) HW Sco Help Me Solve This i Question Help Use the following results from a test for marijuana use, which is provided by a certain drug testing company Among 142 subjects with positive test results, there are 30 false positive results, among 151 negative results there are 5 false negative results. If one of the test subjects is randomly selected, find the probability that the subject tested negative or did not use marijuana (Hint Construct a table.) Use the formal addition rule, where event A is the subject tested negative, event B is the subject did not use marjuana, and P(A and B) denotes the probability that A and B both occur at the same time as an outcome in a trial of a procedure P(A or B)-P(A) P(B)- P(A and B) First, create and till out a table with the values presented in the problem Actually used marijuana? No Yes Positive test result (Test indicated that the subject used marijuana.) Negative test result (Test indicated that the subject did not use marijuana.) 30 Enter your answer in the edit fields and then click Check Answer. parts Clear All Check Answer Close
Answer #1

Total number of subjects = 142 + 151 = 293

Number of people who tested negative = 151

Number of people who did not use marijuana = (142 - 30) + 5 = 117

Number of people who tested negative and did not use marijuana = 151 - 5 = 146

The probability that a randomly selected subject tested negative or did not use marijuana is

= P( subject tested negative) + P(subject did not use marijuana) - P( subject tested negative and did not use marijuana)

151 117 146 293 293 293

151117-146 293

122 293

0.4164 (ans)